# Comparative Studies of Quantum Annealing, Digital Annealing, and Classical Solvers for Reaction Network Pathway Analysis and mRNA Codon Selection

### v1

@@ Abstract
For various optimization problems, the classical time to solution is super-polynomial and intractable to solve with classical bit-based computing hardware to date. Digital and quantum annealers have the potential to identify near-optimal solutions for such optimization problems using a quadratic unconstrained binary optimization (QUBO) problem formulation. This work benchmarks two use cases to evaluate the utility of QUBO solvers for combinatorial optimization problems, in order to determine if a QUBO formulation and annealing-based algorithms have an advantage over classical mixed-integer programming (MIP) and constraint programming (CP) solvers. Various QUBO and solver metrics such as problem mapping, quantitative interconnectivity, penalty structure, solver minimum cost (obtained optimal value) and solver time to solution have been applied to evaluate different QUBO problems. Constrained and unconstrained QUBO solvers are compared including the Fujitsu digital annealer (DA), various D-Wave hybrid quantum annealing solvers (QA, HQA), and the classical MIP/CP solvers HiGHS, Gurobi, SCIP, and CP-SAT. The two industrially relevant use cases are reaction network pathway analysis and mRNA codon selection. For reaction pathway analysis, classical MIP/CP solvers are observed to solve the problem to optimality in reasonable time frames while the DA is not able to do so. For mRNA codon selection, CP-SAT displayed the best performance for standard and large protein datasets (under 1500 amino acids). For the extra-large protein dataset (11000 to 14000 amino acids), the D-Wave Nonlinear HQA solver performed comparably to CP-SAT, outperforming it in minimum cost in 2 out of the 4 problems.
@@ 1.2 Overview of computational solvers
Note that while the DA does not natively support solving HOBO problems, it does provide tools for converting HOBO problems to QUBO problems (that can then be solved) in Fujitsu’s Digital Annealer API [8].
@@ 1.3.3 Quantum annealing
Quantum annealing (QA) is an analog quantum computing based optimization routine. QA aims to solve the SA convergence problem with quantum fluctuations, effectively utilizing tunneling to escape local minima that have high but thin barriers. In QA, the search time grows with $O(e^{\sqrt{N}})$, thus scaling more efficiently than simulated annealing for which the search time on corresponding problems grows with $O(e^{N})$ [16].
@@ 1.4 Commercial and open-source quantum and quantum-inspired annealing solvers
At the time of writing (August 2025), D-Wave offers commercial quantum annealing solvers, such as Advantage and Advantage2, which are superconducting qubit-based quantum annealers with the newest models containing 5000 qubits [15].
@@ 2.2.1 Quantitative connectivity
Density is defined as the ratio of nonzero QUBO terms to the total number of terms ($|x|^{2}$). Such terms are present in a cost or constraint equation $Q$ and in the format of $Q_{ij}x_{i}x_{j}$, with $Q_{ij}\neq 0$ and either $i=j$ or $i\neq j$. A QUBO with all coefficients being zero (all $Q_{ij}=0$) would have a density of 0, and on the other extreme, a QUBO with all coefficients being nonzero (all $Q_{ij}\neq 0$) would have a density of 1. In the context of quantum annealing, density captures the number of couplings between qubits due to the need for embedding onto QPU hardware, and in gate-based quantum computing it represents the number of two-qubit gates needed for the quantum approximation optimization algorithm (QAOA) [37].
@@ 2.2.2 Penalty structure
Penalty separation, or the separation between penalty equation terms (feasibility) and cost function terms (optimality), is also evaluated. If one intends on applying QUBO solvers that do not support constraints terms, penalty terms cannot be expressed in a separate constraint equation and must be manually embedded into the cost function with a scalar (Lagrange multiplier). Such is the case in previous studies applying quantum annealing to ground state energy calculations [40, 6], mRNA codon optimization [34], and reaction network pathway analysis [33]. These problems may require penalty tuning, or an outer optimization loop needed to tune parameters that scale penalty terms in a cost function to help the optimizer balance between focusing on optimality or feasibility [6, 40, 33]. Such a process requires numerous runs of an optimizer and can add significant computational time. In other cases outer optimization may not be needed [34], however the penalty within the cost function can disturb other terms in the cost function and lead to suboptimal solutions.
@@ 2.3 Solver performance metrics
For the D-Wave HQA solvers, different hybrid solver configurations are tested as highlighted in Table 5. The Leap Hybrid Nonlinear (NL) solver which supports decision-variable based problems and constraints, the Leap Hybrid Constrained Quadratic Model (CQM) solver which supports constraints, and the Leap Hybrid Binary Quadratic Model (BQM) solver are benchmarked [41]. Additionally, solvers from the dwave-hybrid open source framework [42] which support embedding constraints into the cost function with a Lagrange multiplier, are tested. The Leap solvers run both CPU and QPU calculations on D-Wave’s cloud infrastructure, whereas for the dwave-hybrid framework the CPU calculations are performed locally. A pure QA solver is not benchmarked due to large problem sizes with heavy interconnectivity that are incompatible with QPU embeddings.
@@ 3.1 Reaction network pathway analysis
DA DA Gurobi HiGHS SCIP CP-SAT
Encoding Unary Log N/A N/A N/A N/A
AC 135.250 132.500 132.417 132.417 132.417 132.417
ATTS [s]* 20.610 30.615 <0.02 <0.02 <0.02 <0.02
@@
DA Gurobi HiGHS SCIP CP-SAT
Encoding Unary N/A N/A N/A N/A
AC 1842.000 455.667 460.333 1170.667 455.667
ATTS [s] 2000.959 3528.728 3530.128 3530.005 3530.501
AC: average cost; average final solution cost over all problems, Gurobi/HiGHS in these cases converge to the optimal (minimum) cost
ATTS: average time to solution
@@ 3.2.1 Benchmarks and results
The Fujitsu DA v4, various D-Wave HQA solvers, CP-SAT (9.14.6206), HiGHS (1.11.0), and SCIP (9.0) were used to solve the QUBO instances. For D-Wave, the following solvers were applied: Leap Hybrid NL Solver (1.22), Leap Hybrid CQM Solver (1.13), Leap Hybrid BQM Solver (2.2), and the dwave-hybrid workflow solvers Kerberos and Parallel Tempering (PT, custom workflow); the dwave-hybrid solvers utilized the Advantage2 quantum computer version 1.5. The classical MIP/CP solvers were run on a server with 48 cores and 98 GB of RAM as were the classical algorithms of the dwave-hybrid Kerberos and PT HQA solvers. The DA and classical MIP/CP solvers solved all problems to optimality, with the D-Wave Leap CQM HQA achieving near-optimal solutions in all cases. A high-level performance comparison is shown in Table 10. Average cost is the average final Hamiltonian cost function value of the solutions found by each solver. Average TTS is the average solve time across all problems, and the standard deviation and maximum of these times are also given. Average QPU usage is the average time the QPU was used by the D-Wave HQA solvers.
@@
Name AC ATTS [s] STTS [s] MTTS [s] AQPU [s]
CP-SAT 188.286 3.062 2.742 8.297 N/A
SCIP 188.286 10.789 14.339 52.513 N/A
HiGHS 188.286 28.476 34.571 113.067 N/A
Digital Annealer 188.286 34.642 4.752 44.17 N/A
Leap NL HQA 188.286 10.957 8.342 30.343 0.493
Leap CQM HQA 188.312 92.888 146.862 500.653 0.077
Leap BQM HQA 269.282 366.639 235.677 700.002 6.410
Kerberos HQA 251.194 361.200 171.657 648.277 0.267
PT HQA 358.833 866.656 787.106 2384.645 0.496
@@
ATTS: average time to solution
STTS: standard deviation of time to solution
MTTS: maximum time to solution
AQPU: average QPU usage
This high-level comparison shows that on the Standard and Large Protein Datasets, the CP-SAT, Leap NL HQA, and DA solvers have the most reliable efficiency in solving to optimality considering average, standard deviation, and maximum time to solution. The Leap CQM HQA is unable to obtain optimal solutions for certain problems although finding near-optimal solutions; the unconstrained D-Wave solvers are unable to identify optimal solutions in a reasonable time. An interesting observation is that the D-Wave proprietary Leap Hybrid BQM solver performs worse than the dwave-hybrid workflow Kerberos solver, and the Leap Hybrid BQM solver uses significantly more QPU time than Kerberos.
The problem scaling behavior as a function of the number of variables illustrates the performance difference between the solvers in more detail, shown in Figure 8. Points that are not filled in represent problems that a solver did not obtain the optimal solution for, and are not included in the lines or curves of best fit. The number of variables in the figure indicates the number of variables for a protein’s QUBO representation, meaning the total number of codons across all amino acids.
@@
As seen in Figure 8 above, the solve time for the CP-SAT, DA, Leap NL HQA, and HiGHS solvers is linear in relation to the problem size, while the solve times for D-Wave and SCIP are exponential. Out of the linear time-scaling solvers, the CP-SAT has the best scaling behavior (lowest slope), followed by the DA, Leap NL HQA, and then HiGHS. The linear time-scaling advantage of these solvers becomes particularly significant for longer, biologically relevant polypeptide sequences. The Leap CQM HQA was not able to identify optimal solutions with problem sizes closer to 4000 variables and above in reasonable amounts of time, however it achieved near-optimal solutions for all problems. The unconstrained D-Wave solvers are unable to identify any optimal solutions in reasonable amounts of time.
@@
The CP-SAT and D-Wave Leap NL HQA solvers, which support the CP formulation of the problem and had the most efficient solve times for the largest problems among the Standard and Large Protein Datasets, were also benchmarked on the Extra-large Protein Dataset. Both solvers were run for 60 seconds and the minimum costs obtained in their solutions are compared. The costs obtained by the solvers for the 4 extra-large proteins are shown in Figure 9. The number of variables in the figure indicates the total number of codons across all amino acids (equivalent to the number of variables for a protein’s QUBO representation, or the total number of options in the CP formulation).
@@
The performance of the CP-SAT and Leap NL HQA solvers on the Extra-large Protein Dataset was comparable, with the CP-SAT outperforming the Leap NL HQA on 2 of the 4 problems, and the Leap NL HQA outperforming CP-SAT on 2 of the 4 problems. For one of the problems in which the Leap NL HQA outperformed CP-SAT, CP-SAT was unable to find a feasible solution in the time limit. Further benchmarking with longer runtimes and with larger problem sizes will be needed to concretely determine which solver performs better for this problem overall.
@@ 4 Lessons learned
This work found CP-SAT and D-Wave Leap NL HQA with a CP formulation to be the most efficient for solving the mRNA codon use case, with the D-Wave NL HQA having comparable performance for the Extra-large Proteins to CP-SAT and solving 2 of these problems closer to optimality than CP-SAT; the DA was the second most efficient in terms of linear scaling for time to solution for the Standard and Large Proteins. The D-Wave Leap CQM HQA was able to solve problems other than the largest ones to optimality but exhibited exponential scaling in time to solution, and the unconstrained D-Wave HQA solvers were unable to identify optimal solutions. For the CRN pathway analysis use case, classical MIP/CP solvers were able to solve the problem to optimality in reasonable runtimes while the DA was not able to. Taking this into account with the above benchmarking framework analysis, the following lessons have been learned:
@@
  • In problems with high density and interconnectivity, when such problems also have high rank-1 dominance classical solvers can potentially solve the problem efficiently with linearizations.
  • CP formulations can offer benefits in runtime compared to MIP and QUBO formulations, for problems which can be more efficiently formulated with decision variables and have rank-1 dominance. This applies to classical and quantum/quantum-inspired solvers supporting decision variables, as seen in the best performing solvers for mRNA codon selection: CP-SAT and D-Wave Leap NL HQA.
  • Quantum and quantum-inspired solvers have overhead costs which can lead to more inefficient solve times for smaller problem sizes, but potentially better scaling behavior to larger sizes compared to classical MIP/CP solvers.
@@
In mRNA secondary structure prediction, conditional value at risk (CVaR)-based variational quantum eigensolver (VQE) run on IBM quantum computers was able to match structure predictions of the classical CPLEX solver in mRNA sequences up to 42 nucleotides; the unconstrained problem formulation for VQE involved embedding a constraint in the cost function. This work demonstrated an exponential scaling in time to solution with classical solvers however did not assess how quantum solvers scale in solving the problem with over 100 qubits. While the work demonstrates progress in quantum solvers for mRNA structure applications, the problem formulation which involved secondary structure base-pairing did not account for tertiary structure contacts [55]. Conversely, in the peptide docking problem, a more complex formulation with 3D distance constraints and steric effects, the authors found that mapping the problem to a HOBO binary optimization matrix was expensive in time, and also had to convert from HOBO to QUBO; these factors contributed to their results highlighting that QUBO optimization is not a good fit for this problem as they benchmarked QUBO-based SA against a constraint-programming solver [36]. Furthermore, for protein folding, a higher-order unconstrained binary optimization (HUBO) formulation was used to model the problem incorporating amino acid interactions (instead of modeling codons and nucleotides like for mRNA problems) and terms up to 5th order; the bias-field digitized counterdiabatic quantum optimization (BF-DCQO) algorithm run on IonQ quantum computers found optimal or high-quality near-optimal solutions in all cases [56].

### v2

@@ Abstract
For various optimization problems, the classical time to solution is super-polynomial and intractable to solve with classical bit-based computing hardware to date. Digital and quantum annealers have the potential to identify near-optimal solutions for such optimization problems using a quadratic unconstrained binary optimization (QUBO) problem formulation. This work benchmarks two use cases to evaluate the utility of QUBO solvers for combinatorial optimization problems, in order to determine if a QUBO formulation and annealing-based algorithms have an advantage over classical mixed-integer programming (MIP) and constraint programming (CP) solvers. Various QUBO and solver metrics such as problem mapping, quantitative interconnectivity, penalty structure, solver minimum cost (obtained optimal value) and solver time to solution have been applied to evaluate different QUBO problems. Constrained and unconstrained QUBO solvers are compared including the Fujitsu digital annealer (DA), various D-Wave hybrid quantum annealing solvers (QA, HQA), and the classical MIP/CP solvers HiGHS, Gurobi, SCIP, and CP-SAT. The two industrially relevant use cases are reaction network pathway analysis and mRNA codon selection. For reaction pathway analysis, classical MIP/CP solvers (especially Gurobi and CP-SAT) are observed to solve the problem to optimality in reasonable time frames. For mRNA codon selection, Gurobi outperformed all other solvers in time to solution for all problem sizes, followed by CP-SAT and the D-Wave Nonlinear (NL) HQA solver.
@@ 1.2 Overview of computational solvers
Note that while the DA does not natively support solving HOBO problems, it does provide tools for converting HOBO problems to QUBO problems (that can then be solved) via Fujitsu’s Digital Annealer API [8].
@@ 1.3.3 Quantum annealing
Quantum annealing (QA) is an analog quantum computing based optimization routine. QA aims to solve the SA convergence problem with quantum fluctuations, effectively utilizing tunneling to escape local minima that have high but thin barriers. In QA, the search time for these certain problems grows with $O(e^{\sqrt{N}})$, thus scaling more efficiently than simulated annealing for which the search time on corresponding problems grows with $O(e^{N})$ [16].
@@ 1.4 Commercial and open-source quantum and quantum-inspired annealing solvers
At the time of writing (August 2025), D-Wave offers commercial quantum annealing solvers, such as Advantage and Advantage2, which are superconducting qubit-based quantum annealers with the newest models containing 5000 qubits [15]. Strong evidence for coherent dynamics in these quantum annealers at short timescales such as 30 nanoseconds has been presented. In practice, many problems require significantly longer anneal times, for which the behavior is expected to depart from coherence and involve thermal effects [21]; tunneling is expected to be mostly incoherent and long-lived entanglement is unlikely.
@@ 2.2.1 Quantitative connectivity
Density is defined as the ratio of nonzero QUBO terms to the total number of terms ($|x|^{2}$). Such terms are present in a cost or constraint equation $Q$ and in the format of $Q_{ij}x_{i}x_{j}$, with $Q_{ij}\neq 0$ and either $i=j$ or $i\neq j$. A QUBO with all coefficients being zero (all $Q_{ij}=0$) would have a density of 0, and on the other extreme, a QUBO with all coefficients being nonzero (all $Q_{ij}\neq 0$) would have a density of 1. In the context of quantum annealing, density captures the number of couplings between qubits due to the need for embedding onto QPU hardware, and in gate-based quantum computing it represents the number of two-qubit gates needed for the quantum approximation optimization algorithm (QAOA) [37, 38].
@@ 2.2.2 Penalty structure
Penalty separation, or the separation between penalty equation terms (feasibility) and cost function terms (optimality), can also be evaluated. If one intends on applying QUBO solvers that do not support constraints terms, penalty terms cannot be expressed in a separate constraint equation and must be manually embedded into the cost function with a scalar (Lagrange multiplier). Such is the case in previous studies applying quantum annealing to ground state energy calculations [41, 6], mRNA codon optimization [34], and reaction network pathway analysis [33]. These problems may require penalty tuning, or an outer optimization loop needed to tune parameters that scale penalty terms in a cost function to help the optimizer balance between focusing on optimality or feasibility [6, 41, 33]. Such a process requires numerous runs of an optimizer and can add significant computational time. In other cases outer optimization may not be needed [34], however the penalty within the cost function can disturb other terms in the cost function and lead to suboptimal solutions.
@@ 2.3 Solver performance metrics
For the D-Wave HQA solvers, different hybrid solver configurations are tested as highlighted in Table 5. The Leap Hybrid Nonlinear (NL) solver which supports decision-variable based problems and constraints, the Leap Hybrid Constrained Quadratic Model (CQM) solver which supports constraints, and the Leap Hybrid Binary Quadratic Model (BQM) solver are benchmarked [42]. Additionally, solvers from the dwave-hybrid open source framework [43] which support embedding constraints into the cost function with a Lagrange multiplier, are tested. The Leap solvers run both CPU and QPU calculations on D-Wave’s cloud infrastructure, whereas for the dwave-hybrid framework the CPU calculations are performed locally. A pure QA solver is not benchmarked due to large problem sizes with heavy interconnectivity that are incompatible with QPU embeddings; such embedding incurs a quadratic overhead [38].
@@ 3.1 Reaction network pathway analysis
Name Encoding AC ATTS*
CP-SAT N/A 132.417 <0.02
Gurobi N/A 132.417 <0.02
HiGHS N/A 132.417 <0.02
SCIP N/A 132.417 <0.02
DA Log 132.500 30.615
DA Unary 135.250 20.610
AC: average cost; average final solution cost over all problems
ATTS [s]: average time to solution in seconds
@@
Name Encoding AC ATTS
Gurobi N/A 455.667 3528.728
CP-SAT N/A 455.667 3530.501
HiGHS N/A 460.333 3530.501
SCIP N/A 1170.667 3530.005
DA Unary 1842.000 2000.959
AC: average cost; average final solution cost over all problems
ATTS [s]: average time to solution in seconds
@@
For genomics optimization problems in general a risk is mapping overhead to QUBO; converting constraints from a problem into a QUBO formulation can incur significant overhead in memory and time. As presented by Maurizio and Mazzola, such overhead is especially common when constraints restrict a large subset of the solution space, very common in genomics problems such as mRNA codon selection [38]; constrained QUBO formulations and solvers can be leveraged to reduce this overhead.
@@ 3.2.1 Benchmarks and results
The Fujitsu DA v4, various D-Wave HQA solvers, Gurobi (12.0.1), CP-SAT (9.14.6206), HiGHS (1.11.0), and SCIP (9.0) were used to solve the QUBO instances. For D-Wave, the following solvers were applied: Leap Hybrid NL Solver (1.22), Leap Hybrid CQM Solver (1.13), Leap Hybrid BQM Solver (2.2), and the dwave-hybrid workflow solvers Kerberos and Parallel Tempering (PT, custom workflow); the dwave-hybrid solvers utilized the Advantage2 quantum computer version 1.5. The classical MIP/CP solvers were run on a server with 48 cores and 98 GB of RAM as were the classical algorithms of the dwave-hybrid Kerberos and PT HQA solvers. The DA and classical MIP/CP solvers solved all problems to optimality, with the D-Wave Leap CQM HQA achieving near-optimal solutions in all cases. A high-level performance comparison is shown in Table 10. Average cost is the average final Hamiltonian cost function value of the solutions found by each solver. Average TTS is the average solve time across all problems, and the standard deviation and maximum of these times are also given. Average QPU usage is the average time the QPU was used by the D-Wave HQA solvers.
@@
Name AC ATTS MTTS STTS AQPU
Gurobi 188.286 0.304 0.682 0.181 N/A
CP-SAT 188.286 3.062 8.297 2.742 N/A
Leap NL HQA 188.286 10.957 30.343 8.342 0.493
SCIP 188.286 10.789 52.513 14.339 N/A
HiGHS 188.286 28.476 113.067 34.571 N/A
DA 188.286 34.642 44.17 4.752 N/A
Leap CQM HQA 188.312 92.888 500.653 146.862 0.077
Kerberos HQA 251.194 361.200 648.277 171.657 0.267
Leap BQM HQA 269.282 366.639 700.002 235.677 6.410
PT HQA 358.833 866.656 2384.645 787.106 0.496
@@
ATTS [s]: average time to solution in seconds
MTTS [s]: maximum time to solution in seconds
STTS [s]: standard deviation of time to solution in seconds
AQPU [s]: average QPU usage in seconds
This high-level comparison shows that on the Standard and Large Protein Datasets, the Gurobi, CP-SAT, and Leap NL HQA solvers have the most reliable efficiency in solving to optimality considering average, maximum, and standard deviation time to solution (TTS). This evaluation focuses on overall performance for the Standard and Large Proteins, hence after average cost, average TTS is considered the most important metric followed by maximum and standard deviation TTS. The Leap CQM HQA is unable to obtain optimal solutions for certain problems although finding near-optimal solutions; the unconstrained D-Wave solvers are unable to identify optimal solutions in a reasonable time. An interesting observation is that the D-Wave proprietary Leap Hybrid BQM solver performs worse than the dwave-hybrid workflow Kerberos solver, and the Leap Hybrid BQM solver uses significantly more QPU time than Kerberos.
The problem scaling behavior as a function of the number of variables illustrates the performance difference between the solvers in more detail, shown in Figure 8. Data points from Table 10 that are not shown represent problems that a solver did not obtain the optimal solution for, and have not been included in obtaining the fitted curves. The number of variables in the figure indicates the number of variables for a protein’s QUBO representation, meaning the total number of codons across all amino acids.
@@
As seen in Figure 8 above, the solve time for the Gurobi, CP-SAT, DA, Leap NL HQA, and HiGHS solvers is linear in relation to the problem size, while the solve times for D-Wave and SCIP are exponential. Out of the linear time-scaling solvers, Gurobi has the best scaling behavior (lowest slope), followed by CP-SAT, DA, Leap NL HQA, and then HiGHS. The linear time-scaling advantage of these solvers becomes particularly significant for longer, biologically relevant polypeptide sequences. The Leap CQM HQA was not able to identify optimal solutions with problem sizes closer to 4000 variables and above in reasonable amounts of time, however it achieved near-optimal solutions for all problems. The unconstrained D-Wave solvers are unable to identify any optimal solutions in reasonable amounts of time.
@@
The Gurobi, CP-SAT, and D-Wave Leap NL HQA solvers, which had the most efficient solve times for the largest problems among the Standard and Large Protein Datasets, were also benchmarked on the Extra-large Protein Dataset. These solvers were run with a time limit of 60 seconds and the minimum costs obtained in their solutions are compared; Gurobi solved all problems to optimality in an average of 14.064 seconds while CP-SAT and the Leap NL HQA solvers were unable to find optimal solutions within the time limit. The costs obtained by the solvers for the 4 extra-large proteins are shown in Figure 9. The number of variables in the figure indicates the total number of codons across all amino acids (equivalent to the number of variables for a protein’s QUBO or MILP representation, or the total number of options in the CP formulation).
@@
Gurobi outperforms the other solvers on the Extra-large Protein Dataset by solving these problems to optimality under the time limit. The performance of the CP-SAT and Leap NL HQA solvers was comparable, with the CP-SAT outperforming the Leap NL HQA on 2 of the 4 problems, and the Leap NL HQA outperforming CP-SAT on 2 of the 4 problems. For one of the problems in which the Leap NL HQA outperformed CP-SAT, CP-SAT was unable to find a feasible solution in the time limit.
@@ 4 Lessons learned
This work found Gurobi with a linearized MILP formulation was the most efficient for solving the mRNA codon use case, with the CP-SAT and D-Wave NL HQA solvers displaying the next best performance. In addition to these solvers, the Fujitsu DA and HiGHS solved mRNA codon selection to optimality with linear scaling in time to solution. The D-Wave Leap CQM HQA was able to solve problems other than the largest ones to optimality but exhibited exponential scaling in time to solution, and the unconstrained D-Wave HQA solvers were unable to identify optimal solutions. For the CRN pathway analysis use case, classical MIP/CP solvers were able to solve the problem to optimality in reasonable runtimes while the Fujitsu DA was not able to.
The Fujitsu DA is designed to solve combinatorial optimization problems using binary variables. In this work we assessed the DA’s capabilities of also encoding integer variables for solving the reaction network pathway analysis use-case. The integer variables encoding contributed to the performance degradation of the Fujitsu DA.
During the course of this study the following lessons have been learned:
@@
  • In problems with high density and interconnectivity, when such problems also have high rank-1 dominance classical MIP/CP solvers can potentially solve the problem efficiently with linearizations.
  • Quantum and quantum-inspired solvers have overhead costs which can lead to more inefficient solve times for smaller problem sizes, but potentially better scaling behavior to larger sizes compared to classical MIP/CP solvers; better scaling behavior of quantum and quantum-inspired solvers was not observed in this work, although the D-Wave Leap NL HQA and DA exhibited linear time scaling for mRNA codon selection.
@@
In mRNA secondary structure prediction, conditional value at risk (CVaR)-based variational quantum eigensolver (VQE) run on IBM quantum computers was able to match structure predictions of the classical CPLEX solver in mRNA sequences up to 42 nucleotides; the unconstrained problem formulation for VQE involved embedding a constraint in the cost function. This work demonstrated an exponential scaling in time to solution with classical solvers however did not assess how quantum solvers scale in solving the problem with over 100 qubits. While the work demonstrates progress in quantum solvers for mRNA structure applications, the problem formulation which involved secondary structure base-pairing did not account for tertiary structure contacts [56]. Conversely, in the peptide docking problem, a more complex formulation with 3D distance constraints and steric effects, the authors found that mapping the problem to a HOBO binary optimization matrix was expensive in time, and also had to convert from HOBO to QUBO; these factors contributed to their results highlighting that QUBO optimization is not a good fit for this problem as they benchmarked QUBO-based SA against a constraint-programming solver [36]. Furthermore, for protein folding, a higher-order unconstrained binary optimization (HUBO) formulation was used to model the problem incorporating amino acid interactions (instead of modeling codons and nucleotides for mRNA problems) and terms up to 5th order; the bias-field digitized counterdiabatic quantum optimization (BF-DCQO) algorithm run on IonQ quantum computers found optimal or high-quality near-optimal solutions in all cases [57], although Tuziemski et al.[58] have highlighted shortcomings in how these results have been benchmarked and that statements about quantum advantage are misleading.